# Mitigating Feature Gap for Adversarial Robustness by Feature Disentanglement

### v1

@@ 5 Conclusion
This paper uncovers a surprising increasing trend in the gap of features between natural and adversarial samples in AFT methods, and further investigates it from the perspective of features. We model features as two feature components, in which the confused features are defined as the latent features leading to the feature gap. Then we propose Adversarial Fine-tuning via Disentanglement (AFD) to bridge the feature gap to enhance robustness. Our approach contains two strategies: disentanglement and alignment. We design the feature disentangler to explicitly separate out the confused features from the features of adversarial samples, while the second strategy aligns the adversarial features with the natural features in the pre-trained model. Experiments demonstrate that AFD effectively mitigates the feature gap and achieves significant robust performance, even surpasses adversarial training methods. Future work will aim at the accuracy-robustness trade-off.
Appendix A Performances of Pre-trained models
Table 7: The natural accuracy of the naturally pre-trained models.
@@ Appendix A Performances of Pre-trained models
We conduct pre-training experiments on multiple datasets and architectures. As shown in Table 7, all of the pre-trained models have much higher accuracies than the fine-tuned models. It indicates that it is reasonable for feature alignment to regard features of the natural samples in the pre-trained models as the ideal unconfused features.
Appendix B Feature Visualization
Figure 6: t-SNE visualization of different methods. The proposed AFD shows a clearer separation in the area of the red box.
To further demonstrate the effectiveness of AFD, Figure 6 visualizes the features of AFT methods using t-SNE (Van der Maaten & Hinton, 2008) on ResNet18 on CIFAR-10. We color each point to match its ground-truth label and highlight focused areas with red boxes. It shows the disentangled features learned by AFD have a much clearer class boundary than those learned with other methods. This indicates that AFD successfully mitigates the confusion between features of distinct classes, leading to a more robust prediction.
Appendix C Environments and Time of Experiments
Table 8: The training time (second) of various methods on ResNet18 on CIFAR10. We show the time for one epoch (${\rm Time_{one}}$) and the expected values of the total time (${\rm Time_{total}}$).
Method Epoch ${\rm Time_{one}}$ ${\rm Time_{total}}$
Standard 100 6.0 590
PGD-AT 120 37 4440
${}^{{\dagger}}$ 4428
TARDES 120 53 6360
${}^{{\dagger}}$ 10 119 1188
${}^{{\dagger}}$ 30 68 2052
${}^{{\dagger}}$ 30 36 1080
vanilla AFT 20 37 1330
RiFT 10 25 4690
AFKD 20 40 1390
ARREST 20 36 1310
AFD 20 46 1510
[…]
As shown in Table 8, TRADES costs the most time and ATAS costs the least time among all the methods. The time cost of RiFT surpasses those of other AFT methods because it needs an adversarially pre-trained model. Our AFD costs less time than TRADES by 4850 seconds and its total training time outperforms that of vanilla AFT by 180 seconds. Considering its advanced robustness, the time cost is worthy and acceptable.

### v2

@@ Conclusion
This paper uncovers a surprising increasing trend in the gap of features between natural and adversarial samples in AFT methods, and further investigates it from the perspective of features. We model features as the intrinsic features and confused features, in which the latter is defined as the specific latent features leading to the feature gap. Then we propose Adversarial Fine-tuning via Disentanglement (AFD) to bridge the feature gap to enhance robustness. We design the feature disentangler to explicitly separate out the confused features from features of adversarial samples. Besides, the disentangled features are aligned with the natural features in the pre-trained model. Experiments demonstrate that AFD effectively mitigates the feature gap and achieves a good trade-off between generalization and robustness.
Acknowledgments
This work was supported in part by the National Natural Science Foundation of China under Grants U22A2096, 62441601, and 62306227, in part by the Shaanxi Province Core Technology Research and Development Project under grant 2024QY2-GJHX-11, in part by the Fundamental Research Funds for the Central Universities under Grants QTZX23042 and ZYTS24142.
Appendix A Further Discussion
Question 1: What are the main differences of the motivation of this paper with the work to explore the existence of robust and non-robust features (Ilyas et al. 2019)?
These papers have different assumptions and definitions. (1) For Ilyas et al. (2019), they assume the features of adversarial samples are composed of robust and non-robust features. Robust features facilitate robustness as well as generalization (i.e., natural accuracy), and non-robust features benefit generalization but damage robustness. (2) In our work, considering the inconsistency of features leads to the vulnerability of DNN, we assume the ideal robust model should have a consistent understanding of natural and adversarial samples, thus extracting similar features from them. Besides, we divide the features into intrinsic and confused features. Intrinsic features contain clean representation. Confused features are induced by adversarial noise and result in a feature gap between natural and adversarial examples, thus causing wrong predictions. The definitions of features are different from Ilyas et al. (2019).
Question 2: As shown in Figure 1 (a), the vanilla AFT also has a low feature distance and its distance is less than the proposed method from epoch 2 to epoch 14. Does it mean that the vanilla AFT gains better robustness?
It is because the vanilla AFT did not converge well in the early stages of training, allowing the corresponding adversarial noise to successfully attack with only a small feature distance. However, it doesn’t change the increasing trend of the feature gap. As shown in Figure 1 (a) after epoch 14, the feature distance of the vanilla AFT grows rapidly and exceeds AFD. Besides, its robust accuracy is lower than AFD all the time (an accuracy disparity of 1.2%-2.8%). When we enlarge the epochs to 40, our AFD rapidly converges to a small value in the feature distance and others still keep a low-speed increasing trend.
@@ Appendix A Further Discussion
Table 6: The clean accuracy of the pre-trained models.
Appendix B Experiments of Pre-trained models
We naturally pre-train all the models with a batch size of 128, training epochs of 100, and a weight decay of ${5.0\times 10^{-4}}$. The learning rate starts at 0.1 and then decays by $\times$0.1 with transition epochs $\{75,90\}$, following Zhang et al. (2019). For AT methods, we train all the models with a batch size of 128, training epochs of 120, and a weight decay of ${2.0\times 10^{-4}}$. The learning rate starts at 0.1 and then decays by $\times$0.1 with transition epochs $\{75,90,100\}$, following Wang et al. (2020).
We conduct pre-training experiments on multiple datasets and architectures. As shown in Table 6, all of the pre-trained models have much higher accuracies than the fine-tuned models. It indicates that it is reasonable for feature alignment to regard features of the natural samples in the pre-trained models as the ideal instrinsic features.
Figure 5: t-SNE visualization of different methods. Our AFD shows a clearer separation in the area of the red box.
Appendix C Feature Visualization
To further demonstrate the effectiveness of AFD, Figure 5 visualizes the features of AFT methods using t-SNE (Van der Maaten and Hinton 2008) on ResNet18 on CIFAR-10. We color each point to match its ground-truth label and highlight focused areas with red boxes. It shows the disentangled features learned by AFD have a much clearer class boundary than those learned with other methods. This indicates that AFD successfully mitigates the confusion between features of distinct classes, leading to a more robust prediction.
Appendix D Supplement
We have conducted experiments on a machine with three NVIDIA GeForce RTX 4090 GPUs. We use a single GPU for ResNet models and three GPUs for WideResNet models. We take three runs for each fine-tuning experiment, and the deviation is 0.43%.
The gap between features of natural and adversarial samples can be observed in all the hidden layers. We select features of the last hidden layer to illustrate the gap since the $l_{p}$ norm of the feature gap in the last hidden layer is the largest.
We have evaluated the feature gap in AFT by measuring the ${L}_{1}$, ${L}_{2}$, and ${L}_{\infty}$ distances between natural and adversarial features. The trends among different metrics are similar to Figure 1 (a).
The results in Figure 3 are obtained from the experiments on ResNet18 on CIFAR-10. The conclusions are similar among distinct datasets and models. The most important of the hyperparameters is $\gamma$, which can be searched in $\{10,25,30,35\}$. Since AFD is insensitive to variations of hyperparameters, it is convenient to apply AFD on various datasets and architectures without searching hyperparameters too many times.